OpenStudy (anonymous):

Solve for w in P = 2w + 2l, if P = 38 and l = 12.

OpenStudy (lgbasallote):

2w + 2L = P 2w = P - 2L w = (p-2l)/2 plug in the values to get w

OpenStudy (anonymous):

P = 2w + 2l P = 38 l = 12 38 = 2w + 2 * 12 38 = 2w + 24 14 = 2w 7 = w

OpenStudy (across):

You are given the equation of the perimeter of a rectangle and are asked to find its width given the perimeter and the length.\[P=2w+2l\implies38=2w+2\cdot12\implies w=7.\]

OpenStudy (anonymous):

thanks guys, you are life savers! :)